# Bézier Splatting for Fast and Differentiable Vector Graphics Rendering

### v3

@@ Abstract
Differentiable vector graphics (VGs) are widely used in image vectorization and vector synthesis, while existing representations are costly to optimize and struggle to achieve high-quality rendering results for high-resolution images. This work introduces a new differentiable VG representation, dubbed Bézier Splatting, that enables fast yet high-fidelity VG rasterization. Bézier Splatting samples 2D Gaussians along Bézier curves, which naturally provide positional gradients at object boundaries. Thanks to the efficient splatting-based differentiable rasterizer, Bézier Splatting achieves 30× and 150× faster per forward and backward rasterization step for open curves compared to DiffVG. Additionally, we introduce an adaptive pruning and densification strategy that dynamically adjusts the spatial distribution of curves to escape local minima, further improving VG quality. Furthermore, our new VG representation supports conversion to standard XML-based SVG format, enhancing interoperability with existing VG tools and pipelines. Experimental results show that Bézier Splatting significantly outperforms existing methods with better visual fidelity and significant optimization speedup. The project page is bezier-splatting.github.io.
Figure 1: This work proposes Bézier Splatting, new differentiable vector graphics (VGs) representation that achieve an order-of-magnitude computational speedup (tested on a NVIDIA RTX 4090 GPU) in comparison with the state-of-the-art differentiable VGs rendering method DiffVG [14].
@@ 1 Introduction
Recently, differentiable VG rasterization gains significant attention, as it allows raster-based algorithms to edit or synthesize VGs through gradient-based optimization. DiffVG [14] is the first differentiable VG framework that leverages the anti-aliasing algorithm to differentiate the vector curves that are inherently discontinuous in pixel space. However, it suffers from slow training and low-fidelity rendering, particularly for high-resolution images. LIVE [17] further introduces a layer-wise coarse-to-fine strategy to improve quality and topology. However, it is highly computationally expensive, requiring 5 hours to vectorize a 2K-resolution image. Additionally, learning-based methods such as Im2Vec [20] train neural networks to map pixels to VGs, but they are restricted to simple graphics and struggle with out-of-domain generalization. Towards scalable applications of VGs, these challenges highlight the need for a differentiable VG method with better efficiency, fidelity, and generalization capability.
This work presents Bézier Splatting, a new differentiable VG representation that optimizes Bézier curves through Gaussian splatting-based rasterization. We sample 2D Gaussian points along Bézier curves and their interior regions, then leverage the Gaussian Splatting framework [11] for efficient curve rasterization. Unlike DiffVG [14] which requires computationally intensive boundary sampling and gradient computation, 2D Gaussians inherently provide direct positional gradients at object boundaries through its differentiable Gaussian formulation, enabling over 150$\times$ faster backward computation over DiffVG for open curves. We further introduce an adaptive pruning and densification strategy that adaptively removes redundant curves while adding new ones to necessary regions during optimization. It helps the optimization process escape the local minima of current spatial distributions of curves, formulating a “global receptive field” for further improving the VG optimization process. As shown in Fig. 1, Bézier Splatting outperforms the state-of-the-art differentiable VG rasterizer DiffVG [14] for both open and closed curves in terms of efficiency and rendering quality. We summarize the contribution of this work as follows.
  • We propose a novel differentiable vector graphic representation, Bézier Splatting, which achieves 30$\times$ faster forward and 150$\times$ faster backward computation, while producing high-quality rendering results.
@@ 2.1 Vectorization and Rasterization
Image vectorization and vector graphic (VG) rasterization are important research topics in computer graphics and computer vision. DiffVG [14] introduces the first differentiable VG rasterization framework, laying a foundation for optimizing and generating vectorized representations through gradient-based methods. It expands the applicability of VGs to a broader range of tasks including image vectorization, text-to-VG generation, and painterly rendering. Based on DiffVG [14], LIVE [17] and O&R [6] further proposes a layer-wise path initialization strategy, vectorizing raster images into compact and semantically consistent VG representations while preserving image topology. Du et al. [3] use linear gradients to fill the regional colors, and Chen et al. [2] propose a specific implicit neural representation to model regional color distributions, enhancing the color representation within closed Bézier curves. However, these DiffVG-based VG rasterization approaches [14, 17, 3, 2] suffer from slow optimization, often requiring several hours to process a 2K-resolution image with 1,024 curves.
With the advancement of deep learning, another line of approaches directly learns deep neural networks for vector synthesis. Lopes et al. [15] combine image-based encoder and VG decoder to generate fonts. Img2Vec [20] integrates the encoder with recurrent neural networks (RNNs) by leveraging sequential modeling for structured vector synthesis. SVGFormer [1] further adopts a Transformer-based architecture [24] to improve the capacity for representing complex geometric structures. More recently, diffusion models [7] have been applied to text-to-VG synthesis [32, 10, 29, 28]. However, existing learning-based VG synthesis approaches are limited to generating simple graphics, and struggle with out-of-domain data. In this work, we propose a novel VG representation dubbed Bézier Splatting that achieves high-fidelity VG rendering within minutes of optimization for high-resolution images.
@@ 2.2 Gaussian Splatting
3D Gaussian Splatting [11, 37] emerges as a promising approach for novel view synthesis (NVS) and attracts significant attention from the community. Its explicit 3D Gaussian-based volumetric representation enables high-fidelity 3D reconstruction, while the differentiable tile-based rasterization pipeline ensures real-time and high-quality rendering. It has been applied to various domains and tasks, such as 4D modeling [16, 27] and 3D scene generation [30, 22]. Several works further improve the Gaussian primitives for better representation quality. SuGaR [4] proposes to approximate 3D Gaussians with 2D Gaussians for enhanced surface reconstruction. 2D Gaussian Splatting [8] directly adopts 2D Gaussians for 3D reconstruction for simplified optimization and improved geometric fidelity. TetSphere splatting[5] further employs tetrahedral meshes as the geometric primitives to achieve high-quality geometry. Particularly for 2D image representation, GaussianImage [34] adopts 2D Gaussians [8] for efficient image representation, achieving a compact and expressive alternative to rasters or implicit representations [21, 18]. Image-GS [35] further enhances it through a content-adaptive compression approach. This work proposes to integrate Gaussian splatting with vector representations for differentiable VG rasterization. By sampling 2D Gaussians on Bézier curves and rasterizing through the efficient Gaussian splatting pipeline, the proposed Bézier Splatting representation enables fast yet high-quality VG rendering, even for high-resolution images with complex structures.
@@ 3.1 Overall
Given a raster image, our goal is to efficiently vectorize it into a VG representation that closely resembles the input while preserving the details. Existing methods, including DiffVG [14] and its following work [17, 2], incur substantial computational costs due to the pixel color accumulation computation. Specifically, DiffVG first constructs a bounding volume hierarchy (BVH) tree to determine the curves that intersect with each individual pixel, then solves equations to precisely determine whether a pixel lies within a region and to compute the inward or outward gradients at boundary points.
To overcome this inefficiency, this work proposes a novel VG representation, Bézier Splatting, which is inspired by the high computational efficiency and expressive fitting capacity of Gaussian splatting [34] for rasterization. Bézier Splatting samples 2D Gaussian points along Bézier curves and their interior regions, then leverages the Gaussian splatting method for efficient rasterization, enjoying the following advantages:
@@
  • The 2D Gaussian representation inherently provides direct position gradients for object boundaries, eliminating the need for additional computations such as boundary sampling and gradient derivation via the Reynolds transport theorem in DiffVG [14];
@@ 3.2 Primitives of Bézier Splatting
To formulate an open curve, we follow DiffVG [14] to adopt three sequentially connected Bézier curves with two control points on each. An open curve requires an additional width parameter to define the stroke thickness. For a closed curve, we adopt two connected Bézier curves, enabling a more efficient sampling in enclosed regions. For closed curves, color filling is applied. The connected Bézier curves in either an open or closed curve share the color parameters, but they have separate opacity parameters to better model the opacity changes along the curves or within closed areas for enriching the texture representation capacity.
2D Gaussians on curves. Our Bézier Splatting novelly associates 2D Gaussians with each Bézier curve. The standard formulation of 2DGS [8] parameterizes each 2D Gaussian by position, color, rotation, scale, opacity, and depth. However, to ensure a compact and differentiable representation, these parameters of Gaussians in our Bézier Splatting are inherited from the corresponding control points. We discuss more details of the sampling of Gaussians, rasterization process, and backaward computation in the following sections.
@@ 3.3 Sampling Gaussians on Bézier Curves
Sampling 2D Gaussians on open curves. We represent an open curve by using $3$ sequential Bézier curves $\mathcal{B}_{i}(t)$ with a degree of $4$ to form a single continuous stroke, by following the same setting as DiffVG [14]. The stroke consists of 10 control points, as the end point of each Bézier curve serves as the start point of the next. To ensure that the final rendering result generates a continuous stroke with consistent width and color, we calculate the x-direction scale of a Gaussian point by the distance between neighboring points by following Eq. 7, while the y-direction scale is a learnable parameter that represents the stroke width.
@@
where $t_{k}\in[0,1]$ are sampled from a normalized cumulative distribution function (CDF). The interpolated control points $\{P^{(k)}_{j}\}$ define intermediate curves that form a dense strip between the two boundaries. Note that this paired structure can naturally extend to cases where each curve is composed of multiple connected Bézier segments, provided that all segments have the same number of control points, similar to the representation used in DiffVG [14].
where $t_{k}\in[0,1]$ are sampled from a normalized cumulative distribution function (CDF). ${P^{(k)}_{j}}$ is interpolated new control points. This non-uniform sampling ensures that points near the boundary curves have small scales, mitigating the influence of interior Gaussians to the exterior of closed area. The interpolated Bézier curves are:
@@
Then, 2D Gaussians are sampled on these interpolated curves, by following the same procedure as open curves (Eq. 3), resulting in a structured and efficient point sampling within the enclosed region. Furthermore, to mitigate artifacts caused by non-convex curve shapes during interpolation, we incorporate the Xing loss from LIVE [17], which enforces convexity constraints on curve shapes, to improve the stability of the interpolation process. The full set of sampled points on the $i$-th Bézier curve is:
@@
3.4 Differentiable Rasterization by Splatting
Once all Gaussians on Bézier curves are sampled, the rasterization process follows the Gaussian splatting pipeline [11], which is very fast in both forward and backward computation. Different from GaussianImage [34], we use $\alpha$-blending for pixel rendering instead of the accumulation-based blending. Accumulation-based blending calculates the pixel value based on all overlapping Gaussians, such that it conflicts with the rendering principles of VGs, where occlusion plays a crucial role in defining vector structures. $\alpha$-blending ensures proper occlusion handling, as it allows foreground elements to contribute more to the final pixel value. The rendering of a pixel is:
@@ 3.4 Differentiable Rasterization by Splatting
Discussion on efficiency. The rasterization process of Bézier Splatting is very efficient, since we directly sample 2D Gaussians from Bézier curves in a differentiable manner, then splat them to the 2D plane. The rasterization pipeline remains end-to-end differentiable while being highly optimized for parallel computation and large-scale matrix operations. As a result, Bézier Splatting is a highly efficient and fully differentiable VG representation, which ensures both fast rendering and gradient-based optimization. It not only preserves the flexibility of VGs but also allows seamless integration into deep learning frameworks, making it well-suited for tasks requiring high-quality and editable vector representations.
@@ 3.5 Optimization
Training objective. Given a raster image $\mathcal{I}\in\mathbb{R}^{H\times W\times 3}$, the goal of image vectorization is to vectorize the image into Bézier curves while ensuring a high-fidelity reconstruction. We first randomly generate a set of Bézier curves to lay on the canvas, then employ the differentiable rasterizer discussed in Sec. 3.4 to render a raster image ^$\hat{\mathcal{I}}\in\mathbb{R}^{H\times W\times 3}$. The Bézier curves can then be optimized through any gradient-based loss functions. In this work, we formulate the optimization objective as minimizing the loss between $\mathcal{I}$ and $\hat{\mathcal{I}}$ while enforcing the curves to be convex. Therefore, we only adopt an $L_{2}$ loss and a Xing loss $L_{\text{Xing}}$ [17], as:
@@
In standard 3D Gaussian Splatting pipeline [11], Gaussians with low opacity or excessive size are pruned, whereas those with high gradient responses are split into two. However, this strategy is not directly applicable to Bézier curves. Note that in volumetric representations, large Gaussians are usually unnecessary for modeling any particular structure. In contrast, VG representations often encompass large uniform regions such as backgrounds or areas with homogeneous colors (e.g., walls). Consequently, the size-based pruning strategy would remove critical structures in VG representations. Similarly, splitting a Bézier curve into two can introduce significant randomness, as high-gradient regions do not always indicate poor reconstruction in VGs. Since VGs assume a uniform color within each enclosed region, complex textures naturally produce high gradients. This does not imply the region should be split, as it may disrupt the semantic consistency of the VG representation.
@@
To address this issue, this work introduces a new pruning and densification strategy to dynamically adjust the density of Bézier curves throughout the optimization process. For pruning redundant Bézier curves, we apply three criteria to ensure a stable and precise optimization process. First, we remove curves with opacity below a dynamic threshold that gradually decreases as optimization progresses, ensuring that weakly contributing curves are eliminated while preserving essential structures. Second, we remove curves with an area below a predefined threshold, as they contribute minimally to the final representation. To remove visually insignificant or noisy Bézier curves, we apply an opacity-based filtering strategy: for per-segment opacities, we discard curves whose middle segment is significantly fainter than both ends—often indicating that a single curve improperly spans a region better represented by multiple curves; otherwise, we discard curves with overall opacity below a threshold (e.g., 0.2) to eliminate globally low-visibility curves. Third, we remove curves that exhibit high color similarity with surrounding curves and have significant overlap, as they provide little additional information and can be pruned without affecting the final representation. By removing such curves, we allow the optimization to introduce more appropriate replacements, leading to a more accurate and visually coherent vectorized representation. For curve densification, we adopt an error-driven curve allocation strategy inspired by LIVE [17]. Specifically, we compute connected error regions, rank them by area, and add new Bézier curves into the highest-error regions. This adaptive redistribution mechanism ensures that curves are allocated to where they would contribute the most to the reconstruction fidelity. The pruning and densification strategy enables a “global receptive field” for redistributing curve density, effectively preserving visual details for high-fidelity rendering while maintaining a compact vector representation.
Table 1: Computational speed for rendering a 2,040$\times$1,344 image with 2,048 curves.
@@
DiffVG [14] Bézier Splatting Speedup DiffVG [14] Bézier Splatting Speedup
Forward pass 141.3ms 4.5ms 31.4$\times$ 85.2ms 14.1ms 6.0$\times$
Backward pass 701.3ms 4.7ms 149.2$\times$ 448.3ms 24.58ms 18.2$\times$
Figure 3: A qualitative comparison of our method and the state-of-the-art differentiable VG rasterization methods, including DiffVG [14], LIVE [17], and LIVSS [26].

### v4

@@ Abstract
Differentiable vector graphics (VGs) are widely used in image vectorization and vector synthesis, while existing representations are costly to optimize and struggle to achieve high-quality rendering results for high-resolution images. This work introduces a new differentiable VG representation, dubbed Bézier Splatting, that enables fast yet high-fidelity VG rasterization. Bézier Splatting samples 2D Gaussians along Bézier curves, which naturally provide positional gradients at object boundaries. Thanks to the efficient splatting-based differentiable rasterizer, Bézier Splatting achieves 30× and 150× faster per forward and backward rasterization step for open curves compared to DiffVG. Additionally, we introduce an adaptive pruning and densification strategy that dynamically adjusts the spatial distribution of curves to escape local minima, further improving VG quality. Furthermore, our new VG representation supports conversion to standard XML-based SVG format, enhancing interoperability with existing VG tools and pipelines. Experimental results show that Bézier Splatting significantly outperforms existing methods with better visual fidelity and significant optimization speedup. The project page is xiliu8006.github.io/Bezier_splatting_project.
Figure 1: This work proposes Bézier Splatting, a new differentiable vector graphics (VGs) renderer that achieves an order-of-magnitude computational speedup in comparison with the state-of-the-art method DiffVG [Li:2020:DVG] (tested on a NVIDIA RTX 4090 GPU).
@@ 1 Introduction
Recently, differentiable VG rasterization gains significant attention, as it allows raster-based algorithms to edit or synthesize VGs through gradient-based optimization. DiffVG [Li:2020:DVG] is the first differentiable VG framework that leverages the anti-aliasing algorithm to differentiate the vector curves that are inherently discontinuous in pixel space. However, it suffers from slow training and low-fidelity rendering, particularly for high-resolution images. LIVE [xu2022live] further introduces a layer-wise coarse-to-fine strategy to improve quality and topology. However, it is highly computationally expensive, requiring 5 hours to vectorize a 2K-resolution image. Additionally, learning-based methods such as Im2Vec [reddy2021im2vec] train neural networks to map pixels to VGs, but they are restricted to simple graphics and struggle with out-of-domain generalization. Towards scalable applications of VGs, these challenges highlight the need for a differentiable VG method with better efficiency, fidelity, and generalization capability.
[…]
  • We propose a novel differentiable vector graphic representation, Bézier Splatting, which achieves an order-of-magnitude computational speedup while producing high-quality rendering results.
@@ 2.1 Vectorization and Rasterization
Image vectorization and vector graphic (VG) rasterization are important research topics in computer graphics and computer vision. DiffVG [Li:2020:DVG] introduces the first differentiable VG rasterization framework, laying a foundation for optimizing and generating vectorized representations through gradient-based methods. It expands the applicability of VGs to a broader range of tasks including image vectorization, text-to-VG generation, and painterly rendering. Based on DiffVG [Li:2020:DVG], LIVE [xu2022live] and O&R [hirschorn2024optimize] further proposes a layer-wise path initialization strategy, vectorizing raster images into compact and semantically consistent VG representations while preserving image topology. Du et al. [Du:2023:IVE] use linear gradients to fill the regional colors, and Chen et al. [Chen_2024_CVPR] propose a specific implicit neural representation to model regional color distributions, enhancing the color representation within closed Bézier curves. However, these DiffVG-based VG rasterization approaches [Li:2020:DVG, xu2022live, Du:2023:IVE, Chen_2024_CVPR] suffer from slow optimization, often requiring several hours to process a 2K-resolution image with 1,024 curves.
With the advancement of deep learning, another line of approaches directly learns deep neural networks for vector synthesis. Lopes et al. [lopes2019learned] combine image-based encoder and VG decoder to generate fonts. Img2Vec [reddy2021im2vec] integrates the encoder with recurrent neural networks (RNNs) by leveraging sequential modeling for structured vector synthesis. SVGFormer [Cao_2023_CVPR] further adopts a Transformer-based architecture [vaswani2017attention] to improve the capacity for representing complex geometric structures. More recently, diffusion models [ho2020denoising] have been applied to text-to-VG synthesis [zhang2024text, jain2023vectorfusion, xing2024svgdreamer, xing2023diffsketcher]. However, existing learning-based VG synthesis approaches are limited to generating simple graphics, and struggle with out-of-domain data. In this work, we propose a novel VG representation dubbed Bézier Splatting that achieves high-fidelity VG rendering within minutes of optimization for high-resolution images.
@@ 2.2 Gaussian Splatting
3D Gaussian Splatting [kerbl3Dgaussians, zwicker2001ewa] emerges as a promising approach for novel view synthesis (NVS) and attracts significant attention from the community. Its explicit 3D Gaussian-based volumetric representation enables high-fidelity 3D reconstruction, while the differentiable tile-based rasterization pipeline ensures real-time and high-quality rendering. It has been applied to various domains and tasks, such as 4D modeling [luiten2023dynamic, Wu_2024_CVPR] and 3D scene generation [yi2023gaussiandreamer, tang2024dreamgaussian]. Several works further improve the Gaussian primitives for better representation quality. SuGaR [guedon2023sugar] proposes to approximate 3D Gaussians with 2D Gaussians for enhanced surface reconstruction. 2D Gaussian Splatting [Huang2DGS2024] directly adopts 2D Gaussians for 3D reconstruction for simplified optimization and improved geometric fidelity. TetSphere splatting[guo2024tetsphere] further employs tetrahedral meshes as the geometric primitives to achieve high-quality geometry. Particularly for 2D image representation, GaussianImage [zhang2024gaussianimage] adopts 2D Gaussians [Huang2DGS2024] for efficient image representation, achieving a compact and expressive alternative to rasters or implicit representations [sitzmann2020implicit, muller2022instant]. Image-GS [zhang2024imagegscontentadaptiveimagerepresentation] further enhances it through a content-adaptive compression approach. This work proposes to integrate Gaussian splatting with vector representations for differentiable VG rasterization. By sampling 2D Gaussians on Bézier curves and rasterizing through the efficient Gaussian splatting pipeline, the proposed Bézier Splatting representation enables fast yet high-quality VG rendering, even for high-resolution images with complex structures.
@@ 3.1 Overall
Given a raster image, our goal is to efficiently vectorize it into a VG representation that closely resembles the input while preserving the details. Existing methods, including DiffVG [Li:2020:DVG] and its following work [xu2022live, Chen_2024_CVPR], incur substantial computational costs due to the pixel color accumulation computation. Specifically, DiffVG first constructs a bounding volume hierarchy (BVH) tree to determine the curves that intersect with each individual pixel, then solves equations to precisely determine whether a pixel lies within a region and to compute the inward or outward gradients at boundary points.
To overcome this inefficiency, this work proposes a novel VG representation, Bézier Splatting, which is inspired by the high computational efficiency and expressive fitting capacity of Gaussian splatting [zhang2024gaussianimage] for rasterization. Bézier Splatting samples 2D Gaussian points along Bézier curves and their interior regions, then leverages the Gaussian splatting method for efficient rasterization, enjoying the following advantages:
@@
  • The 2D Gaussian representation inherently provides direct position gradients for object boundaries, eliminating the need for additional computations such as boundary sampling and gradient derivation via the Reynolds transport theorem in DiffVG [Li:2020:DVG];
@@ 3.2 Primitives of Bézier Splatting
To formulate an open curve, we follow DiffVG [Li:2020:DVG] to adopt three sequentially connected Bézier curves with two control points on each. An open curve requires an additional width parameter to define the stroke thickness. For a closed curve, we adopt two connected Bézier curves, enabling a more efficient sampling in enclosed regions. For closed curves, color filling is applied. The connected Bézier curves in either an open or closed curve share the color parameters, but they have separate opacity parameters to better model the opacity changes along the curves or within closed areas for enriching the texture representation capacity.
2D Gaussians on curves. Our Bézier Splatting novelly associates 2D Gaussians with each Bézier curve. The standard formulation of 2DGS [Huang2DGS2024] parameterizes each 2D Gaussian by position, color, rotation, scale, opacity, and depth. However, to ensure a compact and differentiable representation, these parameters of Gaussians in our Bézier Splatting are inherited from the corresponding control points. We discuss more details of the sampling of Gaussians, rasterization process, and backaward computation in the following sections.
@@ 3.3 Sampling Gaussians on Bézier Curves
Sampling 2D Gaussians on open curves. We represent an open curve by using $3$ sequential Bézier curves $\mathcal{B}_{i}(t)$ with a degree of $4$ to form a single continuous stroke, by following the same setting as DiffVG [Li:2020:DVG]. The stroke consists of 10 control points, as the end point of each Bézier curve serves as the start point of the next. To ensure that the final rendering result generates a continuous stroke with consistent width and color, we calculate the x-direction scale of a Gaussian point by the distance between neighboring points by following Eq. 7, while the y-direction scale is a learnable parameter that represents the stroke width.
@@
where $t_{k}\in[0,1]$ are sampled from a normalized cumulative distribution function (CDF). The interpolated control points $\{P^{(k)}_{j}\}$ define intermediate curves that form a dense strip between the two boundaries. Note that this paired structure can naturally extend to cases where each curve is composed of multiple connected Bézier segments, provided that all segments have the same number of control points, similar to the representation used in DiffVG [Li:2020:DVG].
This non-uniform sampling ensures that points near the boundary curves have small scales, mitigating the influence of interior Gaussians to the exterior of closed area. The interpolated Bézier curves are:
@@
Then, 2D Gaussians are sampled on these interpolated curves, by following the same procedure as open curves (Eq. 3), resulting in a structured and efficient point sampling within the enclosed region. Furthermore, to mitigate artifacts caused by non-convex curve shapes during interpolation, we incorporate the Xing loss from LIVE [xu2022live], which enforces convexity constraints on curve shapes, to improve the stability of the interpolation process. The full set of sampled points on the $i$-th Bézier curve is:
@@
3.4 Splatting-based Differentiable Rasterization
Once all Gaussians on Bézier curves are sampled, the rasterization process follows the Gaussian splatting pipeline [kerbl3Dgaussians], which is very fast in both forward and backward computation. Different from GaussianImage [zhang2024gaussianimage], we use $\alpha$-blending for pixel rendering instead of the accumulation-based blending. Accumulation-based blending calculates the pixel value based on all overlapping Gaussians, such that it conflicts with the rendering principles of VGs, where occlusion plays a crucial role in defining vector structures. $\alpha$-blending ensures proper occlusion handling, as it allows foreground elements to contribute more to the final pixel value. The rendering of a pixel is:
@@ 3.4 Splatting-based Differentiable Rasterization
Discussion on method efficiency. The rasterization process of Bézier Splatting is very efficient, since we directly sample 2D Gaussians from Bézier curves in a differentiable manner, then splat them to the 2D plane. The rasterization pipeline remains end-to-end differentiable while being highly optimized for parallel computation and large-scale matrix operations. As a result, Bézier Splatting is a highly efficient and fully differentiable VG representation, which ensures both fast rendering and gradient-based optimization. It not only preserves the flexibility of VGs but also allows seamless integration into deep learning frameworks, making it well-suited for tasks requiring high-quality and editable vector representations.
@@ 3.5 Optimization
Training objective. Given a raster image $\mathcal{I}\in\mathbb{R}^{H\times W\times 3}$, the goal of image vectorization is to vectorize the image into Bézier curves while ensuring a high-fidelity reconstruction. We first randomly generate a set of Bézier curves to lay on the canvas, then employ the differentiable rasterizer discussed in Sec. 3.4 to render a raster image $\hat{\mathcal{I}}\in\mathbb{R}^{H\times W\times 3}$. The Bézier curves can then be optimized through any gradient-based loss functions. In this work, we formulate the optimization objective as minimizing the loss between $\mathcal{I}$ and $\hat{\mathcal{I}}$ while enforcing the curves to be convex. Therefore, we only adopt an $L_{2}$ loss and a Xing loss $L_{\text{Xing}}$ [xu2022live], as:
@@
In standard 3D Gaussian Splatting pipeline [kerbl3Dgaussians], Gaussians with low opacity or excessive size are pruned, whereas those with high gradient responses are split into two. However, this strategy is not directly applicable to Bézier curves. Note that in volumetric representations, large Gaussians are usually unnecessary for modeling any particular structure. In contrast, VG representations often encompass large uniform regions such as backgrounds or areas with homogeneous colors (e.g., walls). Consequently, the size-based pruning strategy would remove critical structures in VG representations. Similarly, splitting a Bézier curve into two can introduce significant randomness, as high-gradient regions do not always indicate poor reconstruction in VGs. Since VGs assume a uniform color within each enclosed region, complex textures naturally produce high gradients. This does not imply the region should be split, as it may disrupt the semantic consistency of the VG representation.
@@
To address this issue, this work introduces a new pruning and densification strategy to dynamically adjust the density of Bézier curves throughout the optimization process. For pruning redundant Bézier curves, we apply three criteria to ensure a stable and precise optimization process. First, we remove curves with opacity below a dynamic threshold that gradually decreases as optimization progresses, ensuring that weakly contributing curves are eliminated while preserving essential structures. Second, we remove curves with an area below a predefined threshold, as they contribute minimally to the final representation. To remove visually insignificant or noisy Bézier curves, we apply an opacity-based filtering strategy: for per-segment opacities, we discard curves whose middle segment is significantly fainter than both ends—often indicating that a single curve improperly spans a region better represented by multiple curves; otherwise, we discard curves with overall opacity below a threshold (e.g., 0.2) to eliminate globally low-visibility curves. Third, we remove curves that exhibit high color similarity with surrounding curves and have significant overlap, as they provide little additional information and can be pruned without affecting the final representation. By removing such curves, we allow the optimization to introduce more appropriate replacements, leading to a more accurate and visually coherent vectorized representation. For curve densification, we adopt an error-driven curve allocation strategy inspired by LIVE [xu2022live]. Specifically, we compute connected error regions, rank them by area, and add new Bézier curves into the highest-error regions. This adaptive redistribution mechanism ensures that curves are allocated to where they would contribute the most to the reconstruction fidelity. The pruning and densification strategy enables a “global receptive field” for redistributing curve density, effectively preserving visual details for high-fidelity rendering while maintaining a compact vector representation.
[…]
@@
DiffVG [Li:2020:DVG] Bézier Splatting Speedup DiffVG [Li:2020:DVG] Bézier Splatting Speedup
[…]
Figure 3: A qualitative comparison of our method and the state-of-the-art differentiable VG rasterization methods, including DiffVG [Li:2020:DVG], LIVE [xu2022live], and LIVSS [livss].